OpenStudy (koikitty):

During what year will the population of city (a) first exceed that of city (b)? Show all of your work and explain your steps. 2011 2012 2013 2014 156,148 157,788 158,138 160,637 711,195 702,149 696,746 686,674 please dont just give me the answer for salem Y=156,418(1.00891^1) y=157,788(1.00891^2) Y=159,138(1,00891^2) Y=160,937(1,00891^3) for detroit Y=711,195/(1.00891^1) Y= 696,746/(1.00891^1) Y=696,746/(1.00891^2) Y=686,674/(1.00891^3)

OpenStudy (danjs):

You can make a equation for each city, a function of population in terms of time.

OpenStudy (koikitty):

for the equation would it be y=ab^x? @DanJS

OpenStudy (triciaal):

I think a and b are different cities 2 line graphs on the same grid population vs year when Y(a) is > Y(b) what is X , the year notice pop of a is increasing while pop of b is decreasing

OpenStudy (koikitty):

i dont need to graph anything.. so would i use the existing data i have to find out the when the population of cit a would be more the city b? @triciaal

OpenStudy (triciaal):

yes that's what I am saying. what year ?

OpenStudy (koikitty):

okay so i want to find the population of 2016 @triciaal

OpenStudy (triciaal):

I don't know the equations for a or b or how related so I don't know the year

OpenStudy (koikitty):

for salem Y=156,418(1.00891^1) y=157,788(1.00891^2) Y=159,138(1,00891^2) Y=160,937(1,00891^3) for detroit Y=711,195/(1.00891^1) Y= 696,746/(1.00891^1) Y=696,746/(1.00891^2) Y=686,674/(1.00891^3)

OpenStudy (koikitty):

@mathmate can you give me a hand?? just to set up one problem?

OpenStudy (mathmate):

Your idea is sound, but using 1.00891 from Salem is not correct, because Detroit should have a different growth/decay rate. In fact, since the population of Detroit is declining, the rate should be less than 1 and multiplied (instead of divide). Please go through how Salem was set up. You're supposed to learn from it and do it yourself. I have already gone through the first one to serve as an example for the second. I am not allowed to do the work for you, sorry.